an indivdual consumes porduct x and y and spent 500 birr per time period. the prices of the two goods are 60 birr per unit for x and 40 birr per unit for y. the consumers in this case has a utility function expressed as U(x,y)=10xy. A) express the budget equation mathematically


1
Expert's answer
2021-09-20T16:39:44-0400
"60x+40y=500"

"MU_x=10y"

"MU_y=10x"

"\frac{10y}{60}=\frac{10x}{40}"

"x=4"

"y=6"